# Curve fitting on a quantum annealer for an advanced navigation method

### v1

@@ Abstract
We explore the applicability of quantum annealing to the approximation task of curve fitting. To this end, we consider a function that shall approximate a given set of data points and is written as a finite linear combination of standardized functions, e.g., orthogonal polynomials. Consequently, the decision variables subject to optimization are the coefficients of that expansion. Although this task can be accomplished classically, it can also be formulated as a quadratic unconstrained binary optimization problem, which is suited to be solved with quantum annealing. Given the size of the problem stays below a certain threshold, we find that quantum annealing yields comparable results to the classical solution. Regarding a real-word use case, we discuss the problem to find an optimized speed profile for a vessel using the framework of dynamic programming and outline how the aforementioned approximation task can be put into play.
@@ I Introduction
One of the most fundamental approximation tasks is curve fitting, where a function is sought that approximates a given set of data, i.e., the difference between the observed data and the approximation shall become minimal. In this work, we formulate the task of curve fitting as a quadratic unconstrained binary optimization (QUBO) problem [12, 13] that is in principle suited to be solved with quantum computing—particularly quantum annealing [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27].
Furthermore, in this context we discuss the task of finding an optimized speed profile for a vessel in the sense of minimal fuel consumption while arriving just in time using the framework of dynamic programming. Having the voyage of a vessel in mind, it might very well be the case that finding an optimized speed profile requires a lot of support points in order to densly discretize the state space, especially as soon as the latter becomes high-dimensional, and approximations are practically inevitable. Here, we use the QUBO formulation of curve fitting solved on a quantum annealer to approximate the value function with respect to its dependency on the state space variable.
With quantum computing currently being an emerging technology still in its infancy, we do not expect to outperform a conventional computer. Particularly in view of the fact that the only current commercially available quantum annealing devices manufactured by D-Wave [28, 29, 30, 31, 32, 33, 34, 35, 36, 24, 37] are not full-fledged enough yet, and it is a priori not clear whether their use will result in practical advantages for a specific problem [38, 26, 39, 27].
Nevertheless, in our opinion it is worth to explore to what extent quantum annealing can be applied to real-world use cases that arise in an application- and industry-focused environment (see Ref. [27] for an overview) and to judge the quality of the solutions compared to classical methods. The present work lies within that scope.
@@ II Curve fitting as a QUBO
Let $X,Y\subset\mathbb{R}$, $n\in\mathbb{N}$, and $\{(x_{i},y_{i})\}_{i=0}^{n-1}\subset X\times Y$ the observed data points. The task is to find an optimal fit function $f^{\ast}\colon X\to Y$ in the sense that the sum of squares of the differences between the observed data $y_{i}$ and approximated values $f^{\ast}(x_{i})$ is minimized (method of least squares) [40]. With $\mathcal{F}$ denoting a suitable function space, we are concerned with the optimization problem
$\begin{dcases}\min_{f\in\mathcal{F}}Z(f)\,,\\ Z(f)=\sum_{i=0}^{n-1}(y_{i}-f(x_{i}))^{2}\,.\end{dcases}$ (1)
With this problem definition, the best fit is thus given by $f^{\ast}=\operatorname*{arg\,min}_{f\in\mathcal{F}}Z(f)$. For example, if $f(x)=a_{0}+a_{1}x$ with optimization parameters $a_{0},a_{1}\in\mathbb{R}$, the minimization task 1 corresponds to simple linear regression.
@@
$f(x)=\sum_{j=0}^{m-1}c_{j}\phi_{j}(x)\,.$ (2)
For example, such expansions are frequently employed within a certain area of high-energy physics where they aid to handle the numerical complexity of the occurring equations [41, 42]. Two customary choices for $\phi_{j}$ are:
@@
    triangular functions $\Lambda_{j}$ that yield a piecewise-linear approximation. They are defined with respect to supporting points $\tilde{x}_{0}<\dotso<\tilde{x}_{m-1}$, $m>2$, with $\tilde{x}_{0}=x_{0}$ and $\tilde{x}_{m-1}=x_{n-1}$ according to
    $\Lambda_{j}(x)=\begin{dcases}\frac{x-\tilde{x}_{j-1}}{\tilde{x}_{j}-\tilde{x}_{j-1}}&\text{if\leavevmode\nobreak\ }x\in[\tilde{x}_{j-1},\tilde{x}_{j})\,,\\[2.5pt] \frac{\tilde{x}_{j+1}-x}{\tilde{x}_{j+1}-\tilde{x}_{j}}&\text{if\leavevmode\nobreak\ }x\in[\tilde{x}_{j},\tilde{x}_{j+1})\,,\\[2.5pt] 0&\text{otherwise}\,,\end{dcases}$ (4)
    where it is understood that only the first (second) branch defines the nonzero values of the last (first) triangular function $\Lambda_{j=m-1}$ ($\Lambda_{j=0}$).
@@ II.1 Classical solution
vectors $c=[c_{0},\dotsc,c_{m-1}]^{\top}\in\mathbb{R}^{m}$, $b=[b_{0},\dotsc,b_{m-1}]^{\top}\in\mathbb{R}^{m}$, and the symmetric matrix $W\in\mathbb{R}^{m\times m}$ with entries $W_{jk}$, the optimization problem upon inserting Eq. 2 into 1 now reads
@@
The optimal weight coefficients $c^{\ast}$ thus follow from
@@ II.2 Quantum annealing-oriented approach
we have a quadratic function in the components of $c$. With each component of the latter being a real number, it can be written as a difference of two positive real numbers, i.e., $c_{j}=c_{j}^{+}-c_{j}^{-}$ with $c_{j}^{\pm}\in\mathbb{R}_{\geq 0}$. Now, in order to express the optimization problem 7 as a QUBO, whose decision variables’ domain is the binary set $\{0,1\}$, we decompose $c_{j}^{\pm}$ into powers of two, i.e., write them in binary form. This yields the following expression for the expansion coefficients:
$c_{j}=\sum_{r=0}^{L+U}2^{r-L}\hskip 0.83344pt\bigl{(}\psi_{jr}^{+}-\psi_{jr}^{-}\bigr{)}\,,$ (11)
where $L,U\in\mathbb{N}$ are appropriate lower and upper bounds for the binary representation of the expansion coefficients, respectively. $L$ defines the precision with which the decimal places of a real number are represented in a binary expansion. The coefficients $\psi_{jr}^{\pm}\in\{0,1\}$ are then the new decision variables represented by the qubits states of a quantum annealing device.
Here, we tacitly assume that there exists a precision goal that is suitable for all $c_{j}$. In other words, though $L$ and $U$ are (most likely) different for different sets of the underlying to-be-approximated data, they are assumed to be independent of $j$ [^1] This holds provided the values $y_{0},\dotsc,y_{n-1}$ are all of the same order of magnitude. Usually, this is achieved by a proper bijective mapping, i.e., normalization, of the data.
@@
Inserting Eq. 11 into 10 yields an objective function depending on $\psi^{\pm}=[\psi_{00}^{\pm},\psi_{01}^{\pm},\dotsc]^{\top}\in\{0,1\}^{m(L+U+1)}$, i.e., on all coefficients of the binary representation of the expansion coefficients; to wit
$Z(\psi^{\pm})$ $=\sum_{j=0}^{m-1}\sum_{k=0}^{m-1}\sum_{r=0}^{L+U}\sum_{s=0}^{L+U}2^{r+s-2L}W_{jk}$
×$\phantom{=\;}\times\bigl{(}\psi_{jr}^{+}\psi_{ks}^{+}+\psi_{jr}^{-}\psi_{ks}^{-}-\psi_{jr}^{+}\psi_{ks}^{-}-\psi_{jr}^{-}\psi_{ks}^{+}\bigr{)}$
$\phantom{=\;}-\sum_{j=0}^{m-1}\sum_{r=0}^{L+U}2^{r-L+1}b_{j}\hskip 0.83344pt\bigl{(}\psi_{jr}^{+}-\psi_{jr}^{-}\bigr{)}\,.$ (12)
With $d=L+U+1$, this equation can be written as
$Z(\psi^{\pm})$ $=\sum_{p=0}^{md-1}\sum_{q=0}^{md-1}\bigl{(}\psi_{p}^{+}\psi_{q}^{+}+\psi_{p}^{-}\psi_{q}^{-}$
$\phantom{=\;}-\psi_{p}^{+}\psi_{q}^{-}-\psi_{p}^{-}\psi_{q}^{+}\bigr{)}\hskip 0.83344ptW_{pq}^{\prime}$
$\phantom{=\;}-\sum_{p=0}^{md-1}\bigl{(}\psi_{p}^{+}-\psi_{p}^{-}\bigr{)}\hskip 0.83344ptb_{p}^{\prime}\,,$ (13)
@@
where $\psi_{p}^{\pm}=\psi_{\hat{p}\tilde{p}}^{\pm}$, $b_{p}^{\prime}=2^{\tilde{p}-L+1}b_{\hat{p}}$, and $W_{pq}^{\prime}=2^{\tilde{p}+\tilde{q}-2L}W_{\hat{p}\hat{q}}$ with $\hat{p}=\lfloor\hskip 0.83344ptp/d\rfloor$ and $\tilde{p}=p\bmod d$ (analogously for $q$). $b_{p}^{\prime}$ and $W_{pq}^{\prime}$ are the entries of the quantities $b^{\prime}\in\mathbb{R}^{md}$ and $W^{\prime}\in\mathbb{R}^{(md)\times(md)}$, respectively.
Now, in order to cast the objective function 13 into the canonical QUBO form, we introduce the quantities (using block-matrix notation)
$\psi$ $=\biggl{[}\hskip 0.83344pt\begin{matrix}\psi^{+}\\[2.5pt] \psi^{-}\end{matrix}\hskip 0.83344pt\biggr{]}\in\{0,1\}^{2md}\,,$ (16)
$b^{\prime\prime}$ $=\biggl{[}\hskip 0.83344pt\begin{matrix}b^{\prime}\\[2.5pt] -b^{\prime}\end{matrix}\hskip 0.83344pt\biggr{]}\in\mathbb{R}^{2md}\,,$ (19)
@@
which allows us to write the objective function in a rather compact form:
$Z(\psi)=\psi^{\top}W^{\prime\prime}\psi-\psi^{\top}b^{\prime\prime}\,.$ (23)
Furthermore, since $\psi$ is a binary vector, its entries are idempotent, $(\psi_{p}^{\pm})^{2}=\psi_{p}^{\pm}$ for all $p$, and the linear term in Eq. 23 can be written as a contribution that is subtracted from the diagonal elements of $W^{\prime\prime}$. Thus, our final expression for the least-squares optimization problem 1 written as a QUBO reads
$\begin{dcases}\min_{\psi\in\{0,1\}^{2md}}Z(\psi)\,,\\ Z(\psi)=\psi^{\top}Q\hskip 0.83344pt\psi\end{dcases}$ (24)
with $Q=W^{\prime\prime}-\operatorname{diag}(b_{0}^{\prime\prime},b_{1}^{\prime\prime},\dotsc)\in\mathbb{R}^{(2md)\times(2md)}$. Apparently, the number of required qubits depends only on the number of standardized functions $m$ and the precision of the binary decomposition of the expansion coefficients characterized by $L$ and $U$.
III Numerical results
Since we shall compare results obtained on a classical computer using a tabu search [45, 46, 47] with results from a D-Wave quantum annealing device, we briefly sketch how the latter works (see, e.g., Refs. [36, 24, 37, 48] and references therein for more details). In very short: inspired by adiabatic quantum computing [25], quantum annealing as implemented by D-Wave is a heuristic method to find the ground state of an Ising spin model [49, 50]. For a system of spins $s=(s_{0},\dotsc,s_{N-1})\in\{-1,1\}^{N}$, $N\in\mathbb{N}$, the ground state is a spin configuration that globally minimizes the objective function
where $J_{ij}$ and $h_{i}$ denote the nearest-neighbor interaction between the $i\hskip 0.83344pt$th and $j\hskip 0.83344pt$th spin (coupling strength) and an external magnetic field acting on the $i\hskip 0.83344pt$th spin (qubit bias), respectively.
The connection to the corresponding QUBO model with objective function $Z_{\textup{QUBO}}(z)=z^{\top}R\hskip 0.83344ptz$ with $z\in\{0,1\}^{N}$ and an upper-triangular matrix $R\in\mathbb{R}^{N\times N}$ is established by $s_{i}=2z_{i}-1$. The off- and on-diagonal entries of $R$ are mapped onto the coupling strengths and qubit biases, respectively. It is straightforward to show that the Ising and QUBO objective functions differ only by an additive constant, $Z_{\textup{Ising}}(s)=Z_{\textup{QUBO}}(z)+\textup{const.}$., i.e., the solution of a given QUBO problem can be equivalently obtained by finding a ground state of an Ising model.
III.1 Linear regression
Figure 1: Left: data (gray dots) together with optimal regression lines obtained classically ($f_{\textup{cl}}^{\ast}$, solid blue) and by a QUBO solved with tabu search ($f_{\textup{tb}}^{\ast}$, dashed red) as well as quantum annealing ($f_{\textup{qa}}^{\ast}$, dash-dotted green). Right: relative error between the classical and quantum annealing solution.
We proceed with our numerical investigations and begin with the simple yet important case of linear regression. To this end, we choose $n=32$ and consider the mock data ($i=0,\dotsc,n-1$)
$x_{i}=\frac{i}{n-1}\,,\qquad y_{i}=\frac{1}{2}x_{i}+\frac{1}{4}+\textup{noise}\,,$ (26)

### v2

@@ Abstract
We explore the applicability of quantum annealing to the approximation task of curve fitting. To this end, we consider a function that shall approximate a given set of data points and is written as a finite linear combination of standardized functions, e.g., orthogonal polynomials. Consequently, the decision variables subject to optimization are the coefficients of that expansion. Although this task can be accomplished classically, it can also be formulated as a quadratic unconstrained binary optimization problem, which is suited to be solved with quantum annealing. Given the size of the problem stays below a certain threshold, we find that quantum annealing yields comparable results to the classical solution. Regarding a real-world use case, we discuss the problem to find an optimized speed profile for a vessel using the framework of dynamic programming and outline how the aforementioned approximation task can be put into play.
Mathematics Subject Classification (2020): 65D10, 81P68, 90C39
@@ I Introduction
One of the most fundamental approximation tasks is curve fitting, where a function is sought that approximates a given set of data, i.e., the difference between the observed data and the approximation shall become minimal. In this work, we formulate curve fitting as a quadratic unconstrained binary optimization (QUBO) problem [12, 13] that is suited to be solved with quantum computing—particularly quantum annealing [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27].
Furthermore, in this context we discuss the task of finding an optimized speed profile for a vessel in the sense of minimal fuel consumption while arriving just in time using the framework of dynamic programming. Having the voyage of a vessel in mind, it might very well be the case that finding an optimized speed profile requires a lot of support points in order to densely discretize the state space, especially as soon as the latter becomes high-dimensional, and approximations are practically inevitable. Here, we use the QUBO formulation of curve fitting solved on a quantum annealer to approximate the value function with respect to its dependency on the state space variable.
With quantum computing currently being an emerging technology still in its infancy, we do not expect to outperform a conventional computer. Particularly in view of the fact that the only current commercially available quantum annealing devices manufactured by D-Wave [28, 29, 30, 31, 32, 33, 34, 35, 36, 24, 37] are not full-fledged enough yet, and it is a priori not clear whether their use will result in practical advantages for a specific problem [38, 26, 39, 27]. Other application- and industry-focused developement toward quantum annealing devices is carried out by, e.g., the NEC Corporation [40, 41, 42] and Qilimanjaro Quantum Tech [43, 44]. There are also quantum annealing-inspired devices tailored to solve QUBO problems, e.g., the coherent Ising machine [45, 46].
Nevertheless, in our opinion it is worth to explore to what extent quantum annealing can be applied to real-world use cases that arise in an application- and industry-focused environment (see Ref. [27] for an overview) and to judge the quality of the solutions compared to classical methods. This work lies within that scope.
@@ II Curve fitting as a QUBO
Let $X,Y\subset\mathbb{R}$, $n\in\mathbb{N}$, and $\{(x_{i},y_{i})\}_{i=0}^{n-1}\subset X\times Y$ the observed data points. The task is to find an optimal fit function $f^{\ast}\colon X\to Y$ in the sense that the sum of squares of the differences between the observed data $y_{i}$ and approximated values $f^{\ast}(x_{i})$ is minimized (method of least squares) [47]. With $\mathcal{F}$ denoting a suitable function space, e.g., =$\mathcal{F}=\{f\colon\mathbb{R}\to\mathbb{R}\mid f\text{\leavevmode\nobreak\ continuous}\}$, we are concerned with the optimization problem
$\min_{f\in\mathcal{F}}\sum_{i=0}^{n-1}(y_{i}-f(x_{i}))^{2}\,,$ (1)
whose solution is the desired optimal fit $f^{\ast}$. [^1] By construction of minimizing residuals, the least squares method always yields a solution, i.e., an optimal fit $f^{\ast}$ always exists. In our case, where the fit function is written as a finite linear combination of standardized functions according to Eq. 2, the solution is unique if the matrix $W$, defined later in Eq. 6, has full rank. For example, if $f(x)=a_{0}+a_{1}x$, the minimization task 1 corresponds to identifying the optimal parameters $a_{0}^{\ast},a_{1}^{\ast}\in\mathbb{R}$ for an ordinary simple linear regression model.
@@
$f(x)=\sum_{j=0}^{m-1}c_{j}\phi_{j}(x)$ (2)
with $c_{j}\in\mathbb{R}$. Finding an optimal fit $f^{\ast}$ thus translates into finding optimal coefficients $c_{j}^{\ast}$ (for given $\phi_{j}$). For example, such expansions are frequently employed within a certain area of high-energy physics where they aid to handle the numerical complexity of the occurring equations [48, 49]. Two customary choices for $\phi_{j}$ are:
@@
    triangular functions $\Lambda_{j}$ that yield a piecewise-linear approximation. They are defined with respect to supporting points $\tilde{x}_{0}<\dotso<\tilde{x}_{m-1}$, $m\geq 2$, with $\tilde{x}_{0}=x_{0}$ and $\tilde{x}_{m-1}=x_{n-1}$ according to
    $\Lambda_{j}(x)=\begin{dcases}\frac{x-\tilde{x}_{j-1}}{\tilde{x}_{j}-\tilde{x}_{j-1}}&\text{if\leavevmode\nobreak\ }x\in[\tilde{x}_{j-1},\tilde{x}_{j})\\[2.5pt] \frac{\tilde{x}_{j+1}-x}{\tilde{x}_{j+1}-\tilde{x}_{j}}&\text{if\leavevmode\nobreak\ }x\in[\tilde{x}_{j},\tilde{x}_{j+1})\\[2.5pt] 0&\text{otherwise}\end{dcases}$ (4)
    for $j=0,\dotsc,m-1$, where it is understood that only the first (second) branch defines the nonzero values of the last (first) triangular function $\Lambda_{j=m-1}$ ($\Lambda_{j=0}$).
@@ II.1 Classical solution
vectors $c=[c_{0},\dotsc,c_{m-1}]^{\top}\in\mathbb{R}^{m}$, $b=[b_{0},\dotsc,b_{m-1}]^{\top}\in\mathbb{R}^{m}$, and the symmetric matrix $W\in\mathbb{R}^{m\times m}$ with entries $W_{jk}$, the optimization problem upon inserting Eq. 2 into 1 now takes the form
@@
The optimal weight coefficient vector $c^{\ast}$ thus follows from
@@ II.2 Quantum annealing-oriented approach
we have a quadratic function in the components of $c$. In order to express the optimization problem 7 as a QUBO, whose decision variables’ domain is the set $\{0,1\}$, we write each coefficient $c_{j}$ in binary fixed-point format and use two’s complement to represent negative values. This yields the following expression for the coefficients:
$c_{j}=\sum_{r=0}^{d-1}\psi_{jr}\sigma_{r}2^{r-p}\,,$ (11)
where $\psi_{jr}\in\{0,1\}$, $d\in\mathbb{N}$ defines how many binary digits are used, $p=0,\dotsc,d-1$ indicates the location of the fixed decimal point, and
$\sigma_{r}=\begin{cases}-1&\text{if\leavevmode\nobreak\ }r=d-1\\[2.5pt] 1&\text{otherwise}\end{cases}$ (12)
accounts for two’s complement, i.e., the most significant bit enters the sum with a negative sign. For example, for $p=0$, Eq. 11 allows to represent all integers between $-(2^{d-1})$ and $12^{d-1}-$ 1. The quantities $\psi_{jr}$ are then the new decision variables represented by the qubit states of a quantum annealing device.
In principle, one could choose different values for $d$ and $p$ for each coefficient $c_{j}$ matching the individual order of magnitude and location of the decimal point of each coefficient. In practice, however, these are (usually) not known a priori. Thus, and for the sake of simplicity, we choose the same $d$ and $p$ for all coefficients.
@@
Now, inserting Eq. 11 into 10 yields a new objective function depending on $\psi=[\psi_{00},\psi_{01},\dotsc]^{\top}\in\{0,1\}^{md}$, i.e., on all coefficients of the binary representation of the expansion coefficients; to wit
$Z(\psi)$ $=\sum_{j,k=0}^{m-1}\sum_{r,s=0}^{d-1}\psi_{jr}\psi_{ks}\sigma_{r}\sigma_{s}W_{jk}2^{r+s-2p}$
$\phantom{=\;}-\sum_{j=0}^{m-1}\sum_{r=0}^{d-1}\psi_{jr}\sigma_{r}b_{j}2^{r-p+1}\,.$ (13)
This equation can also we written as
$Z(\psi)=\sum_{\mu,\nu=0}^{md-1}\psi_{\mu}W_{\mu\nu}^{\prime}\psi_{\nu}-\sum_{\mu=0}^{md-1}\psi_{\mu}b_{\mu}^{\prime}\,,$ (14)
@@
where $\psi_{\mu}=\psi_{\hat{\mu}\bar{\mu}}$, $b_{\mu}^{\prime}=2^{\bar{\mu}-p+1}\sigma_{\bar{\mu}}b_{\hat{\mu}}$, and $W_{\mu\nu}^{\prime}=2^{\bar{\mu}+\bar{\nu}-2p}\sigma_{\bar{\mu}}\sigma_{\bar{\nu}}W_{\hat{\mu}\hat{\nu}}$ with $\hat{\mu}=\lfloor\mu/d\rfloor$ and $\bar{\mu}=\mu\bmod d$ (analogously for $\nu$). $b_{\mu}^{\prime}$ and $W_{\mu\nu}^{\prime}$ are the entries of the quantities $b^{\prime}\in\mathbb{R}^{md}$ and $W^{\prime}\in\mathbb{R}^{(md)\times(md)}$, respectively. These definitions allow us to write the objective function in a rather compact form:
$Z(\psi)=\psi^{\top}W^{\prime}\psi-\psi^{\top}b^{\prime}\,.$ (15)
Finally, in order to cast this objective function into the canonical QUBO form, we exploit that the entries of the binary vector $\psi$ are idempotent, $\psi_{\mu}^{2}=\psi_{\mu}$ for all $\mu=0,\dotsc,md-1$. The linear term $\psi^{\top}b^{\prime}$ can therefore be written as a contribution that is subtracted from the diagonal elements of $W^{\prime}$. Thus, our final expression for the least-squares optimization problem 1 written as a QUBO reads
$\min_{\psi\in\{0,1\}^{md}}\psi^{\top}Q\psi$ (16)
with the QUBO matrix
$Q=W^{\prime}-\operatorname{diag}(b_{0}^{\prime},\dotsc,b_{md-1}^{\prime})\in\mathbb{R}^{(md)\times(md)}\,.$ (17)
In a last step, the solution $\psi^{\ast}$ of the QUBO problem is plugged into Eq. 11 to obtain the actual, real-valued optimal coefficients $c_{j}^{\ast}=\sum_{r}\psi_{jr}^{\ast}\sigma_{r}2^{r-p}$.
Apparently, the number of required qubits, i.e., the size of the QUBO problem, depends only on the number of standardized functions $m$ used for the expansion of the fit function, see Eq. 2, and the number of digits $d$ used for the binary representation of the expansion coefficients. In other words, the size of the QUBO matrix $Q$ is independent of the number of the to-be-approximated data points $n$ (see Eq. 1); the latter enters only implicitly in the course of the determination of the entries of $Q$ (see Eqs. 5 and 6).
III Numerical experiments
Since we shall compare results obtained on a classical computer using a tabu search [52, 53] with results from a D-Wave quantum annealing device, we sketch very briefly the idea behind the latter (see, e.g., Refs. [36, 24, 37, 54] and references therein for more details). Based on the concept of adiabatic quantum computing [25] and relaxing some of its properties/conditions to facilitate a feasible realization, quantum annealing is a metaheuristic method [24] to find low-energy solutions—including the ground state(s)—of an Ising spin model [55, 56]. For a system of spins $s=[s_{0},\dotsc,s_{N-1}]^{\top}\in\{-1,1\}^{N}$, $N\in\mathbb{N}$, a ground state is a spin configuration that globally minimizes the objective function
where $J_{ij}\in\mathbb{R}$ and $h_{i}\in\mathbb{R}$ denote the nearest-neighbor interaction between the $i\hskip 0.83344pt$th and $j\hskip 0.83344pt$th spin (coupling strength) and an external magnetic field acting on the $i\hskip 0.83344pt$th spin (qubit bias), respectively.
The connection to the corresponding QUBO model with objective function $Z_{\textup{QUBO}}(z)=z^{\top}R\hskip 0.83344ptz$ with $z\in\{0,1\}^{N}$ and an upper-triangular matrix $R\in\mathbb{R}^{N\times N}$ is established by the transformation $s_{i}=2z_{i}-1$ for $i=0,\dotsc,N-1$. The off- and on-diagonal entries of $R$ are appropriately mapped onto the coupling strengths and qubit biases, respectively. It is then straightforward to show that the Ising and QUBO objective functions differ only by an additive constant in the sense that
$Z_{\textup{Ising}}(s)\bigr{\rvert}_{s_{i}=2z_{i}-1}=Z_{\textup{QUBO}}(z)+\textup{const.}\,,$. , (19)
i.e., an approximate or exact solution of a given QUBO problem can be equivalently obtained by finding low-energy or ground states of an Ising model, respectively.
III.1 Sample data
Before we continue with our actual results, we define the sample data $\{(x_{i},y_{i})\}_{i=0}^{n-1}$ (see Section II) that shall be used for our numerical experiments. The abscissa values are evenly spaced over the unit interval according to
$x_{i}=\frac{i}{n-1}\in[0,1]\,.$ (20)
For the ordinates, we choose the following four different types:
@@ III.1 Sample data
[…]